Use the given data to find the minimum sample size required to estimate the population proportion Margin of error 0.008, confidence level: 98%; p and q unknown
Answer by rothauserc(4718)

minimum sample size(n) = N * X/(X +N -1)
:
X = (z*)^2 * p * q/(margin of error)^2
:
margin of error = 0.008
:
confidence level = 98%, therefore z* = 2.33
:
since p and q are not known, we make the following estimates p = 0.50, q = 0.50, population size(N) is 100,000
:
X = (2.33)^2 * (0.50) * (0.50)/(0.008) = 169.6531
:
n = 100000 * 169.6531/(169.6531 +100000 -1) = 169.3675
:
**************************
minimum sample size is 170
**************************
: